What is the magnitude of your displacement when you follow directions that tell you to walk 250 m in one direction, make a 90° turn to the left and walk 330 m, then make a 30° turn to the right and walk 125 m?
1
Expert's answer
2019-10-14T09:15:29-0400
"\bold{d_1}=(250,0)m"

"\bold{d_2}=(0,330)m"

"\bold{d_3}=(125\sin{30},125\cos{30})m"

"\bold{d=d_1+d_2+d_3}=(250+125\sin{30},330+125\cos{30})m"

"\bold{d}=(312.5,438.3)m"

"d=\sqrt{312.5^2+438.3^2}=538\ m"